OpenStudy (anonymous):

Could someone please helpme figure this out? I had to write an equation in slope intercept form, point slope form, or standerd form. Based on the givin information.. Number 1: Slope is 2, Y intercept is -4 Number 2 : is x intercept of 6 and a y intercept of 3 number 3: is passes through (1,2) with slope -5/3 Please explain to me what to do im so lost..

OpenStudy (precal):

number one is using the point slope form y=mx+b you are given your slope m=2 you are given that the y intercept is -4 that means b=-4 just sub those values y=mx+b y=2x-4

OpenStudy (precal):

number two requires algebra

OpenStudy (precal):

we need to find the slope but we know that b=3

OpenStudy (precal):

m=-1/2